Calculate the volume (in mL) of 0.170 M NaOH that
Answer #1

No of mole of MOPS = 0.397*0.0429 = 0.017 mol

pH = pka + log(salt/acid)

7.55 = 7.18+ log(x/(0.017-x))

x= 0.012 mol

No of mole of NaOH added = 0.012 mol

Volume of NaOH = n/M = 0.012 / 0.17 = 0.0706 L

                    = 70.6 ml